suppose a planet has a radius of 5,000 km. starting at the point on its surface at 0 degrees longitude and 0 degrees latitude, find the distance to the point on its surface at 30 degrees longitude and 60 degrees latitude
Answer #1

dlon = lon2 - lon1 = 30 degrees
dlat = lat2 - lat1 = 60 degrees
a = (sin(dlat/2))^2 + cos(lat1) * cos(lat2) * (sin(dlon/2))^2 = 1/4 +0.033= 0.283
c = 2 * atan2( sqrt(a), sqrt(1-a) )
d = R * c (where R is the radius of the Earth)= 5000 x c km